OpenStudy (anonymous):

If there are 15 questions on a test and we are using a 100 point scale, how many points in each question worth? How many do you need to answer correctly to pass the test if a passing score is 70?

OpenStudy (solomonzelman):

x ÷ 15 = each question is worth.

OpenStudy (solomonzelman):

where x=100

OpenStudy (solomonzelman):

each question - 6.67

OpenStudy (solomonzelman):

6.67b = 70 b=70 ÷ 6.67 =10.49 ≈ (rounding up to find the least number of questions to pass) 11

OpenStudy (anonymous):

i need help with two more problem 1.1+5n-1=0 2.3=-5m+4m

OpenStudy (solomonzelman):

1.1+5n-1=0 and 2.3=-5m+4m I think it is written incorrectly, please check

OpenStudy (anonymous):

number one is 1+5n-1=0 number two is 3=-5m+4m

OpenStudy (solomonzelman):

ohh, I thought it is a system of equations.

OpenStudy (solomonzelman):

1+5n-1=0 → 5n=0 → n=0 3=-5m+4m → 3 = -m → m=-3